OpenStudy (anonymous):

The radii of two circles is 3cm and 9cm. find the ratio of the two circumferences. a. 1/9 b. 1/3 c. 2/3

OpenStudy (anonymous):

a. 1/9 becouse\[ratio=\pi r _{1} ^{2}/\pi r _{2} ^{2}=r _{1}^{2}/r _{2}^{2}\]

OpenStudy (anonymous):

Can you help with another one?

OpenStudy (anonymous):

Find the length of the arc in the circle if the radius of the circle is 24 cm. and the degree of the arc is 90. a. 24 pie b. 12 pie c. 6 pie

OpenStudy (anonymous):

hall circle is 360º = 2Pie 1/4 of circle 90º =Pie/2 arc length is r*(angle in radians) so 24*Pie/2=12Pie

OpenStudy (anonymous):

A circle with a diameter of 16 and an area of 64 pie. find the radius. a. 8 b. 4 c. 6

OpenStudy (anonymous):

radius=1/2 diameter so r=8.

OpenStudy (anonymous):

The circle has a radius of 15. find the diameter and area.

OpenStudy (anonymous):

a. D= 30 A= 225pie b. D= 30 A=900 pie c. D= 15 A = 225 pie

OpenStudy (anonymous):

@myko

OpenStudy (anonymous):

You may want to start a new question for that.

OpenStudy (anonymous):

d=r*2=15*2=30 A =Pir^2=Pi15^2=pi225

OpenStudy (anonymous):

a circle has a diameter of 12. find the radius and area.

OpenStudy (anonymous):

dude, all quaestions are the same. Just do like i showed you